OpenStudy (anonymous):

twelve network printers can produce a set of documents in 20 hours. how long would it take 15 printers to get the same job done? (the time is inversly proportional tot the # of printers used for job)

OpenStudy (anonymous):

i have no clue where to start?

OpenStudy (anonymous):

9hrs

OpenStudy (radar):

I think it would be more than that, if they doubled the printers they would half the time that would be 10 hrs and your answer is even less than that. 15/12 : 20/x 15x=12*20 15x=240 x=16 hrs.

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

are you still around?